Does 65 grams of calcium chloride stand for the mass number of calcium chloride ?


1
Expert's answer
2021-02-15T03:19:48-0500

There exist one calcium atom and two chlorine atoms in the structure of calcium chloride.

The molecular mass of this drug is

Ca=1×40=40

Cl=2×35.5=71

CaCl²=40+71= 111 g/mol

Hence, molecular formula of CaCl² = 111 g/mol